OpenStudy (anonymous):

functions f(x) and g(x) are shown below. f(x) = x^2 g(x) = x^2 + 8x + 16 In which direction and by how many units should f(x) be shifted to obtain g(x)? A. Left by 4 units B. Right by 4 units C. Left by 8 units D. Right by 8 units

OpenStudy (anonymous):

A - paste the equations into a google search to see the graphs

OpenStudy (ranga):

f(x) = x^2 g(x) = x^2 + 8x + 16 = (x + 4)^2 If you replace x by (x + a) in f(x) you move left by a units (a > 0) If you replace x by (x - a) in f(x) you move right by a units (a > 0) Here we replace x by (x + 4) to go from f(x) to g(x). Therefore f(x) must be moved ....?

OpenStudy (jdoe0001):

do you know what a "perfect square trinomial" is?

OpenStudy (anonymous):

yeah i know what it is @jdoe0001

OpenStudy (anonymous):

How do I solve this?

OpenStudy (ranga):

f(x) = x^2 g(x) = x^2 + 8x + 16 = (x + 4)^2 We get g(x) from f(x) by replacing x with (x + 4) That means f(x) has been moved left by 4 units.